Evaluate the expression
5P2


Thanks!

Found 2 solutions by venugopalramana, rapaljer:
Answer by venugopalramana(3286)

FORMULA FOR NPR =N!/(N-R)!
5P2=5!/(5-2)!=5!/3!=5*4*3*2*1/3*2*1=5*4=20

Answer by rapaljer(4671)

Easier formula to remember and use:
5P2 means a permutation of 5 things taken 2 at a time. Start with the first number (which in this case is 5), and count down for a total of 2 numbers:
5P2 = 5*4 = 20

No extra charge for a few more:

5P3 means to start with 5 and count down for a total of 3 numbers:
5P3 = 5*4*3 = 60

7P1= 7
7P2 = 7*6 = 42
10P2= 10*9=90
10P3 = 10*9*8 = 720
11P2= 11*10 = 110
6P4 = 6*5*4*3 = 360
etc.

Isn't that simple??

R^2 at SCC